OpenStudy (anonymous):

A moving body undergoes uniform acceleration while travelling in a straight line between points X,Y and Z. The distance XY and YZ are both 40m. The time to travel from X to Y is 12s and from Y to Z is 6.0s. What is the acceleration of the body? a)0.37 b)0.49 c)0.56 d)1.1

OpenStudy (anonymous):

YZ: d=40 t=0.6 v0=? a=? d=v0*t+0.5at^2 40=0.6v0+0.5*a*0.36 =>0.6*v0+0.18*a = 40 v0 = (40-0.18*a)/0.6 XY: v0 = ? d=40 t=12 v=v0 of YZ a=? d = v0*y+0.5a*t^2 but v0 = v-a*t so, d = v-a*t+0.5*a*t^2 40=.... substitute and solve

OpenStudy (anonymous):

Thank you so much! :)